# Boosting Large-scale Parallel Training Efficiency with C4 : A Communication-Driven Approach

### v1

@@ 3. Design
In this section, we will elaborate on how our AI infrastructure tackles the two fundamental challenges identified earlier: ensuring stability and optimizing the performance of parallel training. We will explain the rationale behind our choice of system architecture, the compromises we have accepted, and offer an in-depth analysis of the particular technical hurdles we encountered.
3.1. Stability of Parallel Training
@@
As indicated in Section 2, the GPU cluster’s resource utilization faces challenges due to increasing task sizes and hardware instability. With task scaling being an ongoing trend, we must address hardware failures through two strategies: 1) decreasing hardware failure rates and 2) systematically tolerating such failures. The first strategy is crucial but beyond this paper’s scope, and we briefly mention that temperature control is key (Takeaways-1). Methods like Dynamic Voltage and Frequency Scaling (DVFS) help prevent GPU overheating but can affect performance consistency. We improved cooling with faster fans and better air conditioning, achieving temperature regulation with minimal cost impact. The industry is moving towards more efficient cooling solutions like liquid cooling in future AI infrastructures.
Regarding the second strategy, we have attempted numerous solutions in collaboration with related teams. In general, traditional cloud applications employ online fault tolerance solutions. For instance, they tolerate computational faults by using redundant computations (Poledna, 2007; Mukherjee et al., 2002; Smolens et al., 2006), provide highly reliable storage through erasure coding (EC) (Weatherspoon and Kubiatowicz, 2002; Xia et al., 2015; Dimakis et al., 2006; Aguilera et al., 2005) and/or triple-replica techniques, and endure network anomalies through multi-path (dual uplinks) strategies (Han et al., 2006; Lu et al., 2018). However, for high-performance computing applications, the current mainstream approach is to use offline fault tolerance methods (Koo and Toueg, 1987). For example, applications periodically save checkpoints, and when a system failure occurs, the tasks are restarted from the last checkpoint. Of course, all computations between the last checkpoint and the time of the failure will be discarded. For our large-scale parallel training tasks, we have adopted a relatively pragmatic and hybrid technical strategy.
In general, we utilize proven cloud storage technologies to ensure reliable data storage, while also fully recognizing the separate challenges that computation and networking can present. We had an in-depth discussion with our customers and got a key message, i.e, the underlying system should not introduce any uncertainty into the model training process (Takeaways-2). Therefore, we have only two options for fault-tolerant computing: either perform online redundant computations or prepare backup resources to replace failed nodes. Given the high cost of GPUs, it’s more cost-effective to use backup resources rather than online redundancy for fault tolerance. Consequently, we’ve allocated 64 backup GPUs across 8 servers for every 1024 GPUs on 128 servers, ensuring identical communication and performance for parallel training on any 128 servers from this 136-server pool.
Regarding networking, the industry typically adopts single-port cards such as 1*400G to diminish the chances of hash collisions. Nonetheless, this configuration introduces reliability concerns, as a port malfunction can force tasks to restart from the most recent checkpoint. Our methodology, which reflects the insights gained from customer feedback, emphasizes swift failure detection and the maintenance of network connectivity. This strategy enables higher-layer applications to react more adeptly, such as scheduling timely restarts or initiating on-demand checkpoints that are strategically based on recent checkpointing activities. Pursuing this direction, we have opted for 2*200G dual uplinks to enhance reliability, while simultaneously addressing network hash collisions to sustain application performance. In essence, we believe that ensuring maximum reliability at each layer, coupled with cross-layer optimization, is vital for achieving the most efficient and reliable overall system (Takeaways-3).
Our approach: C4D. In summary, our system’s fault-tolerant architecture includes reliable data storage through erasure coding, network reliability via dual uplinks and multi-path communication, and handling of computational faults through checkpoints and redundancy. The core of this system is the C4D (C4 Diagnose) subsystem, designed to quickly detect hardware failures for prompt task restarts, as shown in Fig. 3.
Figure 3. The overview of the fault-tolerant system.
@@
C4D leverages two key observations (Takeaways-4): (1) parallel training tasks exhibit a regular and predictable running rhythm, enabling us to precisely identify abnormal behaviors; (2) the Bulk Synchronous Parallel (BSP) execution model in parallel training requires regular synchronization points. And these synchronization points are used as anchors for measuring anomalies.
Based on above observations, we developed the C4D to facilitate online fault detection. Fig. 4 illustrates the key elements of the C4D, which includes an enhanced Collective Communication Library (CCL), a central C4D master, and the C4a (C4 agent) that acts as an intermediary. We opt to use enhanced CCL because state-of-the-art CCLs, such as NCCL (Jeaugey, 2017), lack the capability for online monitoring of collective communication operations.
Figure 4. The architecture of C4D.
@@
C4D Monitoring. Fig. 5 shows our enhanced CCL with four layers, similar to contemporary CCLs. We’ve expanded the bottom three layers for monitoring functionalities. The communicator layer tracks communicator IDs, rank counts, and rank assignments. The operation layer monitors collective operation types, algorithms, data types, element counts, and operation durations and counts. The transport layer gathers connection specifics like RDMA IP and QP numbers, and message statistics including counts, sizes, and durations of transfers. It’s not trivial to collect above information at runtime, with low cost and high accuracy. To precisely monitor communication kernel execution patterns, we refined all related CDUA kernels to log their start and completion times directly, as CPU timestamps and CUDA Events are ineffective or inaccurate for this purpose.
Figure 5. The Enhanced CCL.
@@
C4D analysis. Based on above gathered informations, we can detect four common types of errors that frequently occur in our clusters, i.e., communication hang, non-communication hang, communication slow, and non-communication slow. Detecting the first two error types is relatively easy and will not be discussed in-depth here, and we will focus on the more complex task of identifying instances of the slow syndrome.
@@
Case 1: Communication Slow Detection. Before delving into the details, we will give a brief introduction to the data flow in collective communication. Considering the allreduce operation in data parallel phase as an example, where workers need to average gradients across all model replicas. Frameworks and CCLs will split messages into chunks and submitting them to the network sequentially. Since all workers adhere to the same data splitting rules, ensuring that message sizes remain consistent when posted to the network. Therefore, we can identify communication slowdowns at the transport layer by monitoring and comparing the latency of messages in different workers.
An illustrative example is shown in Fig. 6. Communication delays across workers are mapped into a two-dimensional matrix, where each element represents the delay between a pair of workers, identified by their y-coordinate (source) and x-coordinate (destination). High values in the matrix help identify slow connections: a single high-value point indicates a specific connection bottleneck, a row of high values suggests a problem with the source, and a column of high values points to an issue with the destination.
Figure 6. The symptom for communication slow.
@@
Case 2: Non-Communication Slow Detection. Consider the ring-based algorithm in allreduce operation as an example. In this configuration, all participating workers are connected sequentially, forming a ring-like structure. Each worker only communicate with its immediate neighbors in the sequence, specifically the “previous rank” and the “next rank”. Particularly, workers receive a data chunk from the “previous rank”, perform a reduction operation with their local data chunk, and then pass the resultant data to the “next rank”. In fact, due to the requirement that the receiver must first prepare the receive buffer and notify the sender before data transmission can occur, there is an implicit ”receiver-driven” scheduling logic. Therefore, we can diagnose non-communication slows, such as those caused by computation or data loading, by looking at how long the receiver waits for data. For instance, if the receiver experiences a non-communication slow, the sender is likely already waiting, allowing for quick data receipt once scheduled. Conversely, if the sender has a non-communication slow, it won’t send data promptly even after receiving the receiver’s scheduling signal, resulting in lengthy wait times for the receiver.
3.2. The Scalability of Parallel Training
@@
Parallel training performance depends on the efficiency of single-node computation, data access, and collective communication. While single-node efficiency can be improved through optimizations like mixed precision (NVIDIA, [n. d.]d) and transformer engine (NVIDIA, [n. d.]b), and data access efficiency through caching mechanisms like Alluxio (Li, 2018) , this paper focuses on collective communication efficiency, a key factor in training scalability.
@@
The first issue is not the focus of this paper, so for the sake of completeness, we provide a brief introduction here. Once the data volume for one-to-one communication is fixed, its consumption of network resources is directly proportional to the distance the data travels through the network. To reduce the data transmission distance, we employ two optimization strategies. First, we minimize the network diameter through innovative network architecture. AI training servers internally provide high-speed NVLINK interconnects, which are inherently part of the network (Takeways-5). Thus, our network is actually a hierarchical topology, with NVLINK forming the first layer and the RDMA network across different servers comprising the second layer. Furthermore, we have utilized dual-uplink technology, which not only improves network reliability but also allows for an increase in the radix of switch chips. Clearly, for a given number of endpoints and a specified network topology, the greater the switch radix, the smaller the network diameter (Kim, 2008). Second, we utilize network topology-aware scheduling techniques to ensure that the two ranks needing to communicate are as close as possible within the network.
Based on the above optimizations, we can achieve good performance in most cases. For instance, for single tasks with a small scale, all communications can be completed within a single layer of leaf switches. But for training tasks of a larger scale or in scenarios where multiple tasks coexist, further optimization is essential. The underlying problem in both of the aforementioned scenarios is actually the same: there is a significant amount of traffic that requires forwarding by the spine switches, resulting in a high volume of traffic collisions.
@@
Our approach: C4P. To address this challenge, we introduced the C4P (C4 Performance) system, engineered to reduce unnecessary communication costs. C4P optimizes communication for concurrent jobs and link failures by: (1) identifying and avoiding faulty links at task start-up, (2) balancing RDMA QPs, i.e., connections, across paths for load distribution, and (3) dynamically adapting QP workloads in response to network changes and traffic conflicts. Essentially, C4P is a traffic engineering technique that aims to minimize network congestion by regulating the transmission paths of each data flow within the network. This concept isn’t new, but it’s not commonly employed in traditional data centers, because of the massive connections in network. C4P is practical in parallel training scenarios mainly because the traffic characteristics generated by parallel training are significantly different from those of conventional cloud applications. That is, parallel training tasks involve a small number of data flows but transmit large volumes of data. This presence of a few elephant flows makes it feasible for C4P to meticulously plan the route for each data flow (Takeways-6).
C4P follows C4D’s software structure but with key differences, as shown in Fig. 7. Firstly, the C4P master acts as a control center for multiple jobs or tenants, unlike the single-job-focused C4D master. Additionally, C4P’s CCL can request path allocations for communicating workers, while C4D’s collects monitoring data. Lastly, C4P’s master allocates communication paths, whereas C4D’s master handles fault detection and diagnosis.
@@
Similar to prior studies (Dong et al., 2021), we use path-probing for refined traffic engineering. C4P first isolates and discards malfunctioning links between leaf and spine switches, creating a healthy-link network. The C4P master performs full-mesh path probing via randomly selected servers per leaf switch, identifying and cataloging reliable paths. On connections setup, the CCL prompts path requests to the C4P master, which responses selected path by specifying the source ports of RDMA connections. The master ensures traffic from the same NIC is balanced between left and right ports by forbidding the paths from left ports to right, and vice versa. Additionally, traffic from servers under the same leaf switch is distributed over all available spine switches to prevent network congestion. The CCL constantly evaluates message completion times on various paths and prioritizes the fastest for data transfer. If the optimal QP’s queue is full, the next best is chosen. This approach helps maintain low transmission delays despite path errors or traffic congestion.
Figure 7. The working flow of C4P.
The deployment approach for C4P parallels C4D’s, with both systems embedded into user images to facilitate integration with Kubernetes (K8s). A significant distinction, however, is that the C4P master operates on a system-wide level, offering global access, in contrast to the C4D master, which serves a more localized role, restricted to individual job scopes. Details beyond this will not be reiterated here to prevent redundancy.
4. Evaluations
4.1. Setup
@@
As shown in Table 2, we outline the system configurations of our operational clusters, from which we have gathered the evaluation data for C4D. The assessments pertaining to C4P, on the other hand, were carried out on a subset of this cluster. To prevent any interference with other ongoing jobs in the cluster and to ensure the integrity of our evaluation results, we have allocated a specific segment of the cluster as a controlled testing environment. This testbed comprises 16 nodes, equipped with a total of 128 GPUs. And all the nodes are directly connected to 8 dedicated leaf switches, ensuring an exclusive environment for our testing activities.
@@
The nodes within this high-performance cluster are outfitted with 8 NVIDIA H800 GPUs and 8 BlueField-3 NICs (NVIDIA, [n. d.]e). Each NIC provides two physical 200Gbps ports, which are bonded to create a single logical 400Gbps port. These NICs are integrated into a 3-Tier Clos network (Clos, 1953), configured in a Fat-Tree topology with a 1:1 oversubscription rate, ensuring optimal performance and bandwidth distribution. Leveraging Broadcom’s Trident4 (BROADCOM, [n. d.]b) as leaf and Tomahawk4 as spine switches (BROADCOM, [n. d.]a), the cluster is capable of supporting over 10,000 GPUs. In a single pod, which constitutes a two-tier subnet, it can accommodate 512 GPUs, showcasing its expansive scalability.
Table 2. Configurations
@@
Frameworks Megatron-LM (Shoeybi et al., 2020), DeepSpeed (Rajbhandari et al., 2020)
@@
The effectiveness of C4D is evaluated with one of our real-world LLM training workloads. And the model for this job contains 175 billion parameters. Undertaking distributed training with 2400 GPUs, the model typically requires over a month for a complete training cycle from scratch. In contrast, the efficiency of C4P is assessed using both collective communication benchmarks and three representative LLM training jobs. To ensure unbiased evaluation results for the communication benchmarks, we configured them to utilize a ring-based algorithm. The specific configurations of the models involved in the testing, such as parallel strategy and zero-optimizer level, will be detailed alongside the presentation of the evaluation outcomes.
4.2. Results
4.2.1. The Evaluations on C4D’s Effectiveness.
We have reviewed the historical logs of the jobs from one of our inner customers, including their start and terminated timestamps, and whether a job is succeeded or failed by some reasons. Our primary emphasis has been placed on a representative job that demands 2400 GPUs and necessitates over a month to complete its model training process. As shown in Table 3, we compare the error-induced downtime of this job before (Jun, 2023) and after (Dec, 2023) the proposed fault-tolerant system was deployed. The data clearly indicates that the deployment of C4D led to a substantial decrease in downtime, with an approximately 30x reduction, from 31.19% down to just 1.16%.
Table 3. Error-induced Downtime Statistics

### v2

@@ III Design
To tackle the key issues identified earlier, i.e., the overheads induced by local failures and unnecessary slowdowns in communication and other processes. We will explain the rationale behind our choice of system architecture, the compromises we have accepted, and offer an in-depth analysis of the particular technical hurdles we encountered.
III-A Mitigating Error-induced Downtime
@@
As demonstrated in the previous section, the significant hardware instability in computing nodes is a primary factor contributing to the inefficiency of AI clusters. To mitigate the impact of hardware failures, we have adopted a two-fold approach: 1) reducing hardware failure rates and 2) systematically tolerating such failures. While the first aspect is beyond the scope of this paper, we briefly mention the importance of temperature control. Techniques such as Dynamic Voltage and Frequency Scaling (DVFS) have been utilized to prevent GPU overheating, although they can impact performance consistency. We have enhanced cooling through increased fan speed and improved air conditioning, allowing for temperature regulation with minimal cost impact. The industry is also moving towards more efficient cooling solutions such as liquid cooling for future AI infrastructures.
With respect to the second aspect, we have collaborated with relevant teams to explore multiple solutions. Traditional cloud applications typically utilize online fault tolerance techniques. These may involve withstanding computational faults through redundant computations [41, 32, 50], ensuring reliable storage with erasure coding (EC) [55, 56, 11, 1] and/or triple-replica techniques, and mitigating network anomalies with multi-path (dual uplinks) strategies [18, 29]. For high-performance computing applications, the prevailing approach involves offline fault tolerance methods [24]. This includes the periodic saving of checkpoints, allowing tasks to be restarted from the last checkpoint in the event of system failure. Yet, any computations made between the last checkpoint and the time of the failure will be lost. In our large-scale parallel training systems, we have implemented a practical and hybrid technical strategy. This includes adopting proven online fault-tolerant technologies in distributed storage and networking systems. While for computing resources, it’s more cost-effective to use backup resources and offline fault-tolerant techniques, given the pricing of GPU products. As a result, we have allocated 64 backup GPUs across 8 servers for every 1024 GPUs on 128 servers, ensuring consistent communication and performance for parallel training on any of the 128 servers within this 136-server pool.
Our approach: C4D. The system that we designed to handle computational faults is shown in Fig. 4. The core of this system is the C4D (C4 Diagnose) subsystem, designed to quickly detect hardware failures for prompt task restarts. Initially, C4 agents monitor the operational status of training workers and transmit the data to a centralized master. The master then evaluates the well-being of the training workers by comparing the gathered information. If any irregularities are detected, it informs the job steering service to isolate the problematic nodes and restart the job from the most recent valid checkpoint. Simultaneously, it also sends the events to a background root cause analysis system for offline diagnosis.
Figure 4: The working flow of C4D.
@@
The effectiveness of C4D is grounded in two key insights: (1) parallel training tasks exhibit a regular and predictable running rhythm, enabling us to precisely identify abnormal behaviors; (2) the Bulk Synchronous Parallel (BSP) execution model in parallel training requires regular synchronization points. And these synchronization points are used as anchors for measuring anomalies. Building on these insights, we developed the C4D to facilitate online fault detection, as shown in Fig. 5. The key elements of the C4D includes the enhanced communication library ACCL, a central C4D master, and the C4a (C4 agent) that acts as an intermediary. Here, ACCL is enhanced to provide the capability for online monitoring of collective communication operations.
Figure 5: The architecture of C4D.
@@
C4D Monitoring. Fig. 6 shows the architecture and enhancement in ACCL. We have extended the monitoring capabilities of the bottom three layers, namely the communicator, operation, and transport layers, in a top-down fashion. We gather information on communicators, encompassing communicator IDs, involved devices, and device ranks. Additionally, we record details about collective operations, such as operation types, communication algorithms, data types, element counts, and duration. Furthermore, we track the startup and completion of specific collective operations and assign each operation a sequence. Lastly, we collect data in the transport layer, including connection information (e.g., source and destination IPs, QP numbers, and source ports), as well as message details like message counts, sizes, and transfer duration. Notably, it’s not trivial to collect above information at runtime, with low cost and high accuracy. To precisely monitor communication kernel execution patterns, we refined all related CDUA kernels to log their start and completion times directly, as CPU timestamps and CUDA events are ineffective or inaccurate for this purpose.
Figure 6: The Enhancement of ACCL.
@@
C4D analysis. By leveraging the gathered information from ACCL, we can build the automatic error-detection system to deal with the four common types of errors that frequently occur in our clusters, i.e., communication hang, non-communication hang, communication slow, and non-communication slow. Detecting the first two error types is relatively easy and will not be discussed in-depth here, and we will focus on the more complex task of identifying instances of the slow syndrome.
@@
Communication Slow Detection: Before delving into the details, we will give a brief introduction to the data flow in collective communication. Considering the allreduce operation in data parallel phase as an example, where workers need to average gradients across all model replicas. Frameworks and CCLs will split messages into chunks and submitting them to the network sequentially. Since all workers adhere to the same data splitting rules, ensuring that message sizes remain consistent when posted to the network. Therefore, we can identify communication slowdowns at the transport layer by monitoring and comparing the completion time of messages in different workers. An illustrative example is shown in Fig. 7. Communication delays across workers are mapped into a two-dimensional matrix, where each element represents the delay between a pair of workers, identified by their y-coordinate (source) and x-coordinate (destination). Large numbers in the matrix help identify slow connections: a single large number indicates a specific connection bottleneck, a row of large numbers suggests a problem with the source, and a column of large numbers points to an issue with the destination.
Figure 7: The symptom for communication slow.
@@
Non-Communication Slow Detection: Consider the ring-based algorithm in allreduce operation as an example. In this configuration, all participating workers are connected sequentially, forming a ring-like structure. Each worker only communicate with its immediate neighbors in the sequence, specifically the “previous rank” and the “next rank”. Particularly, workers receive a data chunk from the “previous rank”, perform a reduction operation with their local data chunk, and then pass the resultant data to the “next rank”. In fact, due to the requirement that the receiver must first prepare the receive buffer and notify the sender before data transmission can occur, there is an implicit ”receiver-driven” scheduling logic. Hence, successful communication completion in each step depends on both the sender and receiver being in a ready state. This implicit dependency is then spread to following receivers, forming a chain of pending receivers. By comparing the wait time of receivers, we can pinpoint the ranks that are experiencing non-communication slows, such as those caused by additional computation or data loading costs.
III-B Mitigating Communication Cost
@@
Parallel training performance depends on the efficiency of single-node computation, data access, and collective communication. While single-node efficiency can be improved through optimizations like mixed precision [38] and transformer engine [36], and data access efficiency through caching mechanisms like Alluxio [27] , this paper focuses on collective communication efficiency, a key factor in training scalability.
@@
The first issue is not the focus of this paper. But for the sake of completeness, we provide a brief introduction here. Once the data volume for one-to-one communication is fixed, its consumption of network resources is directly proportional to the distance it travels through the network. To reduce the data transmission distance, we employ two optimization strategies. First, we minimize the network diameter by leveraging high-speed NVLink interconnects, which are inherently part of the network. Second, we utilize topology-aware scheduling techniques to ensure that the two ranks needing to communicate are as close as possible within the network. These optimizations mitigate the probabilities of traffic collision, and are effective for small-scale jobs.
Furthermore, we have also integrated intra-job traffic engineering techniques [13, 14] to prevent traffic collisions. However, in the case of larger AI clusters with multiple concurrent jobs, there is a lack of coordination among independent jobs. Consequently, additional optimization is required to deal with inter-job traffic collision.
Figure 8: The architecture of C4P.
@@
Our approach: C4P. To address this problem, we introduce the C4P (C4 Performance) system, engineered to reduce unnecessary communication costs. Essentially, C4P is a cluster-scale traffic engineering technique that aims to minimize network congestion by regulating the transmission paths of each data flow within the network. This concept isn’t new, but it’s not commonly employed in traditional data centers, because of the massive connections in network. C4P is practical in parallel training scenarios mainly because the traffic characteristics generated by parallel training are significantly different from those of conventional cloud applications. That is, parallel training tasks involve a small number of data flows but transmit large volumes of data. This presence of a few elephant flows makes it feasible for C4P to meticulously plan the route for each data flow.
The software architecture of C4P follows that of C4D but with key differences, as shown in Fig. 8. Firstly, the C4P master acts as a control center for multiple jobs or tenants, unlike the single-job-focused C4D master. Additionally, the C4P master records the numbers of allocated connections on each path, and allocates path for new connections considering the occupied network resources. At last, the CCL is enhanced to support issuing path allocation requests for communicating workers and set the source port accordingly. The key functionalities of C4P can be summarized as: (1) identifying and avoiding faulty links at task start-up, (2) balancing RDMA QPs, i.e., connections, across healthy paths for load distribution, and (3) dynamically adapting QP workloads in response to network changes and traffic conflicts.
@@
The detailed working flow is as follows: In line with previous art [14], we utilizes path-probing to support global traffic engineering. Using this method, we can identify the source ports that will direct traffic along specific paths and verify the integrity of those paths. Initially, C4P identifies and eliminates faulty links between leaf and spine switches to establish a sub-network composed of healthy links. To do this, the C4P master performs full-mesh path probing via randomly selected servers per leaf switch, identifying and cataloging the healthy paths. Upon establishing connections, ACCL submits path allocation requests to the C4P master, which replies with the source ports of RDMA connections. The master ensures traffic from the same NIC is balanced between left and right ports by forbidding the paths from left ports to right, and vice versa. Additionally, traffic from servers under the same leaf switch is distributed over all available spine switches to prevent network congestion. The ACCL constantly evaluates message completion times on various paths and prioritizes the fastest for data transfer. This approach helps maintain low transmission delays despite path errors or traffic congestion.
IV Evaluations
IV-A Setup
@@
As shown in Table II, we outline the system configurations of our operational clusters, from which we have gathered the evaluation data for C4D. The assessments pertaining to C4P, on the other hand, were carried out on a subset of this cluster. To prevent any interference with other ongoing jobs in the cluster and to ensure the integrity of our evaluation results, we have allocated a specific segment of the cluster as a controlled testing environment. This testbed comprises 16 nodes, equipped with a total of 128 GPUs. And all the nodes are directly connected to 8 dedicated leaf switches, ensuring an exclusive environment for our testing activities.
@@
The nodes within this high-performance cluster are outfitted with 8 NVIDIA H800 GPUs and 8 BlueField-3 NICs [39]. Each NIC provides two physical 200Gbps ports, which are bonded to create a single logical 400Gbps port. These NICs are integrated into a 3-Tier Clos network [7], configured in a Fat-Tree topology with a 1:1 oversubscription rate. Leveraging Broadcom’s Trident4 [4] as leaf and Tomahawk4 as spine switches [3], the cluster is capable of supporting over 10,000 GPUs. In a single pod, which constitutes a two-tier subnet, it can accommodate 512 GPUs.
TABLE II: Configurations
@@
Frameworks Megatron-LM [49], DeepSpeed [45]
@@
The effectiveness of C4D is evaluated with one of our real-world LLM training workloads. And the model for this job contains 175 billion parameters. Training with 2400 GPUs, it requires more than a month for a complete training cycle from scratch. In contrast, the efficiency of C4P is assessed using both collective communication benchmarks and three representative LLM training jobs. To ensure unbiased evaluation results for the communication benchmarks, we configured them to utilize a ring-based algorithm. The specific configurations of the models involved in the testing, such as parallel strategy and zero-optimizer level, will be detailed alongside the presentation of the evaluation outcomes.
IV-B Results
IV-B1 The Evaluations on C4D’s Effectiveness.
We have reviewed the historical logs of the jobs from one of our inner customers, including their start and terminated timestamps, and whether a job is succeeded or failed by some reasons. Our primary emphasis has been placed on a representative job that demands 2400 GPUs and necessitates over a month to complete. As shown in Table III, we compare the error-induced downtime of this job before (Jun, 2023) and after (Dec, 2023) the proposed fault-tolerant system was deployed. The data clearly indicates that the deployment of C4D led to a substantial decrease in downtime, with an approximately 30x reduction, from 31.19% down to just 1.16%.
TABLE III: Error-induced Downtime Statistics